2. Determine your weight (e.g., the force of gravity on you).
3. Determine what normal force would be required to accelerate you upwards from ground level at a rate of 2.0 or 3.0 m/s^2 (choose one acceleration value).
4. What would the normal force be if the acceleration described in (3) was directed downwards?
1
Expert's answer
2020-10-29T07:01:27-0400

2)


"W=mg=(80)(9.8)=784\ N"

3)


"N=m(a+g)=80(2+9.8)=944\ N"

4)


"N=m(-a+g)=80(-2+9.8)=624\ N"